# ECLARE: Efficient cross-planar learning for anisotropic resolution enhancement

### v1

@@ 4.3 People with MS: T2-w FLAIR simulations
For the 3D-MR-MS dataset, we used the automated white matter lesion segmentation algorithm SELF (Zhang et al., 2024) to produce labels for the GT and super-resolved volumes, comparing results with cDSC. We also included signal recovery metrics PSNR and SSIM as with the previous dataset.
Figure 8: A $4\|1.2$ mid-sagittal slice for a representative subject from the 3D-MR-MS dataset. The first row shows the image domain with the automated lesion segmentation overlaid in neon-green. The second row shows the log Fourier magnitude of the corresponding slice. The third row shows the absolute difference between the log Fourier magnitudes of the SR method and the ground truth. The PSNR, SSIM, and cDSC for the entire image volume of this representative are overlaid in the image domains.
We performed similar analyses for the 3D-MR-MS dataset, but with SynthSR omitted due to the image being T2-w FLAIR. Qualitative results are shown in Fig. 8, where the first row in the image domain overlays the automated segmentation by SELF in neon-green. This T2-w FLAIR is a prime example of the “over-sharpening” artifact from SMORE, such as the hyperintense region just inferior to the ventricle. Many such examples of hyperintense regions just inferior to a dark surface are visible in this result, whereas ECLARE does not exhibit this. Furthermore, ECLARE better recovers from aliasing artifacts induced by the gap, since this image is $4\|1.2$. Again, from the third row, it is apparent that single-contrast INR changes DC.
@@
In Fig. 9, we show the PSNR and SSIM for signal recovery and cDSC for lesion segmentation consistency. Like in the OASIS-3 dataset results, SMORE is the only competitor quantitatively at $3\|0$, though in this case it is not a clear winner. The upper quartile appears to outperform ECLARE, but the mean is lower. These results are consistent across all three metrics. After $3\|0$, ECLARE outperforms the rest.
4.4 Real-world data
We evaluated ECLARE alongside bicubic interpolation, single-contrast INR, and SMORE for the SR of real-world data described in Sec. 4.
Figure 10: Results on real data for which there is no ground truth. The first column shows the nearest-neighbor interpolated LR data and the second column shows the cubic B-spline interpolated result. The third column shows Single-contrast INR, the fourth column shows SMORE, and the final column shows ECLARE. The first row is the ULF scan, the second is SWI, the third is the mouse scan, and the fourth is the spine.
The results of each of these are shown in Fig. 10, with the nearest neighbor interpolation shown as LR, and each method in subsequent columns. We make two quick notes about single-contrast INR here. First, single-contrast INR was unable to process the SWI image because of its memory requirements, so its result is omitted from the figure. Second, single-contrast INR does not include a model of slice thickness in its formulation; the PSF is assumed to be Dirac. This makes single-contrast INR more of an “inpainting” technique during training rather than super-resolution, since at inference-time the blurred signal will not be deblurred, which is detrimental to the results of single-contrast INR.
The advantage of SMORE and ECLARE is apparent by inspection: aliasing is reduced and images have less blurring than bicubic interpolation. This is evident throughout the representative slices on the ULF, Mouse, and Spine images. The zoomed inset for all images is also illustrative of differences between the proposed methods. For the ULF scan, single-contrast INR is unable to estimate an image that is as realistic as both SMORE and ECLARE.
On the ULF and spine images, we notice both SMORE and ECLARE exhibit a kind of periodic banding spanning the image from left to right. On close examination, this band is approximately equivalent to the interpolated signal at that location. We hypothesize that at large $r$ values, it is difficult for either method to fill in the unknown information. However, both SMORE and ECLARE are able to respect the original signal at those locations (likely due to how they were trained) and thus the banding appears. Solving this issue is left to future work.
@@ 5 Discussion and conclusions
In this work, we presented ECLARE as a novel self-SR algorithm for anisotropic volumes that requires no external training data. We presented a model of the imaging process in 2D multi-slice MRI and used ESPRESO to estimate the slice selection profile. With an estimated slice profile, slice thickness and slice separation are both modeled, allowing for training data simulation from the HR in-plane data already present in the anisotropic volume. ECLARE uses a neural network architecture based on WDSR to efficiently train weights from scratch in a short amount of time, taking approximately 5 minutes on a single NVIDIA V100 GPU with a memory footprint under 2 GB. The result is a SR technique that respects the LR image contrast and pathology presence, does not require external training data, and produces HR image volumes with the correct digital resolution in a reasonable amount of time. When compared to four other approaches on two simulated datasets, our approach achieved superior quantitative and qualitative results with respect to signal recovery metrics PSNR and SSIM as well as a downstream task metric—cDSC for brain region and MS lesion segmentation. When looking at results for real-world data, qualitative differences show the utility of ECLARE.
We note two primary limitations of our work here. First, the results from ECLARE do not appear to recover the full isotropic spatial resolution compared to the ground truth images. As apparent from Figs. 4 and 8, high frequency information (details) are absent from the predicted result. The reason for this is two-fold. First, our optimization objective function is solely mean squared error, which is known to lead to blurrier results when used in SR tasks. We also optimized L1, but found nearly identical results qualitatively and quantitatively. Methods that model a distribution, such as adversarial approaches or denoising diffusion probabilistic models, yield improved realness and may lead to improvements in ECLARE. However, the perception-distortion tradeoff (Blau and Michaeli, 2018) asks us to choose between realistic appearances and signal recovery, and we chose the latter. Additionally, modeling distributions potentially introduces the risk of domain dependence. Second, because ECLARE is a self-SR method the in-plane data used as HR reference is not truly HR; it is subject to partial volume effects due to through-plane blurring. This affects ECLARE twice: 1) the target HR data is not representative of true HR slices, and 2) if ESPRESO is used to estimate $h$ but the volume is highly anisotropic, the resulting kernel may be a spike. These issues may be amended by including a reference HR image volume as a single example to learn from.
@@
The SR of medical images has been a research area for decades. Data-driven approaches now dominate the field of SR, yet the problem is far from solved. When the test-time image does not belong to the training domain, performance of trained approaches is undefined. In medical imaging, domain issues may occur across a number of axes, including resolution difference, anatomical variance, presence of pathology, differences in contrast or scan acquisition parameters, noise level and distribution mismatch, or others. When the test-time image does belong to the training domain, it is unknown whether the super-resolved image is reliable for downstream analysis. Further work is needed to investigate the extent to which an SR algorithm is reliable. Analysis is also needed to determine what slice thicknesses, separations, and profile shapes are best suited for SR and reliability. As these problems are addressed in the future, both the field of SR and the ability of SR to make meaningful clinical impact and research impact will advance forward.

### v2

@@ 4.3 People with MS: T2-w FLAIR simulations
One prominent characteristic of MS is the presence of pathology as white matter hyperintense lesions. Data-driven SR algorithms conventionally must include relevant pathology in their training data to be expected to perform well on pathological data. ECLARE, SMORE, and single-contrast INR accomplish this intrinsically since they train on in-plane slices, many of which contain lesions. For the 3D-MR-MS dataset, we used the automated white matter lesion segmentation algorithm SELF [61] to produce labels for the GT and super-resolved volumes, comparing results with cDSC. We also included signal recovery metrics PSNR and SSIM as with the previous dataset.
Figure 8: A $4\|1.2$ mid-sagittal slice for a representative subject from the 3D-MR-MS dataset. Each column shows the MR image with the automated lesion segmentation overlaid in neon-green. The PSNR, SSIM, and cDSC for the entire image volume of this representative are overlaid.
We performed similar analyses for the 3D-MR-MS dataset, but with SynthSR omitted due to the image being T2-w FLAIR. Qualitative results are shown in Fig. 8, where the automated segmentation by SELF is overlaid in neon-green. This T2-w FLAIR clearly shows the “over-sharpening” artifact from SMORE in the hyperintense region inferior to the ventricle. Many such examples of hyperintense regions just inferior to a dark surface are visible in the SMORE result, whereas the ECLARE result does not exhibit this. Furthermore, ECLARE better recovers from aliasing artifacts induced by the gap, since this image is $4\|1.2$.
@@
In Fig. 9, we show the PSNR and SSIM for signal recovery and cDSC for lesion segmentation consistency. Like in the OASIS-3 dataset results, SMORE is best at $3\|0$, though in this case it is not a clear winner. The upper quartile appears to outperform ECLARE, but the mean is lower. These results are consistent across all three metrics. At all resolutions worse than $3\|0$, ECLARE has the best performance among all algorithms.
@@ 5 Discussion and conclusions
In this paper we presented ECLARE, a novel self-SR algorithm for anisotropic volumes that trains solely on data extracted from in-plane slices and thus requires no external training data. We presented a model of the imaging process in 2D multi-slice MRI and used ESPRESO [30] to estimate the slice selection profile. With an estimated slice profile, slice thickness and slice separation are both modeled, allowing for training data simulation from the HR in-plane data already present in the anisotropic volume. ECLARE uses a neural network architecture based on WDSR to efficiently train weights from scratch in a short amount of time, taking approximately 5 minutes on a single NVIDIA V100 GPU with a memory footprint under 2 GB. The result is a SR technique that respects the LR image contrast and presence of pathology, does not require external training data, and produces HR image volumes with the correct digital resolution and FOV. When compared to four other approaches on two simulated datasets, our approach achieved superior quantitative and qualitative results with respect to signal recovery metrics PSNR and SSIM as well as a downstream task metric—cDSC for brain region and MS lesion segmentation. This is true for all investigated resolutions except for $3\|0$, where SMORE performs best. This is likely due to two reasons: 1) SMORE does not have a model of slice gap, and therefore will perform best when no gap is present in the data; and 2) SMORE relies on pre-trained weights and was primarily tested and evaluated using data with resolutions $2\|0$ and $3\|0$.
We note that computation time is a primary limitation of ECLARE. ECLARE must fit a model for each multi-slice 2D MR volume. Although its training time is relatively fast compared to other self-SR methods, it is slower than pre-trained models that perform inference only. It would be possible to pre-train-then-fine-tune ECLARE as SMORE, but this introduces a tradeoff between time and relying on pre-training as a prior more than the in-plane data itself. Future work will investigate these tradeoffs in detail.
@@
The SR of medical images has been a research area for decades. Data-driven approaches now dominate the field of SR, yet the problem is far from solved. When the test-time image does not belong to the training domain, performance of trained approaches is unpredictable. In medical imaging, domain issues may occur across a number of axes, including resolution differences, anatomical variations, presence of pathology, differences in contrast or scan acquisition parameters, noise level and distribution mismatch, or others. Our proposed method ECLARE considers all of these by training on in-plane data and leveraging a forward model inspired by the MR acquisition physics. Because it does not rely on external data for prior information, ECLARE has the potential to be used on any volumetric image and is not limited to scenarios for which a large amount of data—matched in both contrast and resolution—can be used for training.